OpenStudy (anonymous):

double integration of xy/(x^2+y^2)

OpenStudy (anonymous):

The limits for both dy and dx are from 0 to 1. xy*sqty(x^2 + y^2) dy dx Also having a tough time with this one... integrate with with the boundaries y = sqrt(x), y = 2, x = 0. sin(y^3) dy dx

terenzreignz (terenzreignz):

@meghagoyal I don't think double integration makes sense without limits...so are they really 0 to 1 for both x and y as stated above^ ?

OpenStudy (anonymous):

\[\int\limits_{0}^{1}\int\limits_{0}^{1}\frac{xy}{x^2+y^2}dx dy =\int\limits_{0}^{1}\frac{1}{2}[\int\limits_{0}^{1}\frac{2x}{x^2+y^2}dx] y dy=\int\limits_{0}^{1}Log[x^2+y^2] |^1_0 y dy\]

OpenStudy (anonymous):

simplify the second integral by the substitution x^2+y^2 = t